(a) A nurse has two solutions that contain different concentrations of a certain medication. One is a 12.5% concentration and the other is a 5% concentration. How many cubic centimeters of each should she mix to obtain 20 cubic centimeters of an 8% concentration?

Answer by richwmiller(17219)

a+b=20,
0.125*a+0.05*b=0.08*20
a=20-b
0.125*(20-b)+0.05*b=1.6
2.5-0.125b+0.05*b=1.6
-0.075*b=-0.9
b=12.0
a=20-b
a=8 cc at 12.5% b=12 cc at 5%
check
0.125*8+0.05*12=0.08*20
1.0+0.6=1.6
1.6=1.6
ok